2. Solve the IVP (x + x = 3uſt-27) |x(O)=0, x(O)= 2

Solution steps plz
Answer #1

- 2118 2: z + x = 34(2-21 Taking Laplace transform są - sx(0) - x0) +ă =) ł (8 41) = 2+ e 24 3 8 -{u 200303] -2RS s -2118 2